OpenStudy (anonymous):

Find dy/dx of y = (x^3-2x)^ln x?

OpenStudy (anonymous):

\[\left(x^3-2 x\right)^{\log (x)} \left(\frac{\log \left(x^3-2 x\right)}{x}+\frac{\left(3 x^2-2\right) \log (x)}{x^3-2 x}\right) \]

OpenStudy (nikvist):

How did you get it?

OpenStudy (nikvist):

Full answer